OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 29°, b = 26, c = 28

OpenStudy (espex):

\[\frac{b}{\sin{B}} = \frac{c}{\sin{C}}\]